What is the partial pressure of each gas if the following amounts of substances are placed in a 25.0 L container at 25 °C? What is the total pressure of each mixture?

a. 1.570 mol of CH4 and 0.870 mol of CO2
b. 2.63 g of CO and 1.24 g of NO2
1
Expert's answer
2021-10-15T02:58:38-0400

a. Using ideal gas law, find the pressure of CH4 and CO2:


"P_1=\frac{n_1RT}{V}=\frac{1.57·8.314·(25+273)}{25·10^{-3}}=155670\text{ Pa}, \\\space\\
P_2=\frac{n_2RT}{V}=\frac{0.87·8.314·(25+273)}{25·10^{-3}}=86263\text{ Pa},"

the total pressure can be found by Dalton's law:


"P=P_1+P_2=241933\text{ Pa}."


b. Repeat the procedures from (a):


"P_1=\frac{n_1RT}{V}=\frac{2.63·8.314·(25+273)}{25·10^{-3}}=260772\text{ Pa}, \\\space\\
P_2=\frac{n_2RT}{V}=\frac{1.24·8.314·(25+273)}{25·10^{-3}}=122949\text{ Pa}, \\\space\\
P=P_1+P_2=383721\text{ Pa}."